40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  
|
Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking
|
Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking
|
Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking 
|
Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking
|
Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking
|
Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB   
|
Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions
|
Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions
|
Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions
|
Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling  
|
Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling
|
Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise
|
Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed
|
Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed
|
Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey 
|
Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling 
|
Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out
|
Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out
|
Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate 
|
Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises
|
Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises
|
Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled  
|
Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled
|
Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled
|
Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF 
|
|
Message ID: 12979
Entry time: Wed May 10 01:56:06 2017
In reply to: 12976
Reply to this: 12981
12983
|
Author: |
gautam |
Type: |
Update |
Category: |
General |
Subject: |
MICH NB - OL coupling |
|
|
Last night, I tried to estimate the contribution of OL feedback signal to the MICH length error signal.
In order to do so, I took a swept sine measurement with a few points between 50 Hz and 500 Hz. The transfer function between C1:LSC-MICH_OUT_DQ and the Oplev Servo Output point (e.g. C1:SUS-BS_OL_PIT_OUT etc) was measured. I played around with the excitation amplitude till I got coherence > 0.9 for the TF measurement, while making sure I wasn't driving the Oplev error point too hard that side-lobes began to show up in the MICH control signal spectrum.
The Oplev control signal is not DQ-ed. So I locked the DRMI again and downloaded the 16k data "live" for ~5min stretch using cdsutils.getdata on the workstation. The Oplev error point is DQ-ed at 2k, but I found that the excitation amplitude needed for good SNR at the error point drove the servo to the limiter value of 2000cts - so I decided to use the control signal instead. Knowing the transfer function from the Oplev *_OUT* channel to C1:LSC-MICH_IN1_DQ, I backed out the coupling - the transfer function was only measured between 50 Hz and 500 Hz, and no extrapolation is done, so the estimation is only really valid in this range, which looks like where it is important anyways (see Attachment #2, contributions from ITMX, ITMY and BS PIT and YAW servos added in quadrature).
I was also looking at the Oplev servo shapes and noticed that they are different for the ITMs and the BS (Attachment #1). Specifically, for the ITM Oplevs, an "ELP15" is used to do the roll-off while an "ELP35" is employed in the BS servo (though an ELP35 also exists in the ITM Oplev filter banks). I got lost in an elog search for when these were tuned, but I guess the principles outlined in this elog still hold and can serve as a guideline for Oplev loop tweaking.
Coil driver noise estimation to follow
Quote: |
I think the most important next two items to budget are the optical lever noise, and the coil driver noise. The coil driver noise is dominated at the moment by the DAC noise since we're operating with the dewhitening filters turned off.
|
GV 10 May 12:30pm: I've uploaded another copy of the NB (Attachment #3) with the contributions from the ITMs and BS separated. Looks like below 100Hz, the BS coupling dominates, while the hump/plateau around 350Hz is coming from ITMX. |
|
|
|
|
|